Total Production 1999 2000 Price Hotdogs T-Shirts Factories Total Production 1000 2000 10 Price 2 4 100 Total Production 2000 4000 20 100 26. The chart above shows total production in an economy and the prices of each good produced in 1999 and 2000. Let 1999 be the base year, and let the market basket for the CPI be 1 hotdog and 1 T-shirt. Which of the the following statements is true? a. The inflation rate calculated by using the CPI is lower than the inflation rate calculated by using the GDP deflator The inflation rate calculated by using the CPI is higher than the inflation rate calculated by using the GDP deflator The inflation rate calculated by using the CPI is equal to the inflation rate calculated by using the GDP deflator More information is required to compare the inflation rates calculated by using the CPI vs. using the GDP deflator b. c. d.
Answer #1

Option (b).

First we compute CPI and GDP deflator as follows.

(i) CPI

Cost of basket, 1999 = 1 x 1 + 1 x 1 = 1 + 1 = 2

Cost of basket, 2000 = 2 x 1 + 4 x 1 = 2 + 4 = 6

CPI, 1999 = 100 (since 1999 is base year)

CPI, 2000 = (Cost of basket, 2000 / Cost of basket, 1999) x 100 = (6/2) x 100 = 300

Inflation rate = % Change in CPI = (300/100) - 1 = 3 - 1 = 2 = 200%

(ii) GDP Deflator

Nominal GDP = \sum(Current year price x Current year quantity)

1999 = 1 x 1000 + 1 x 2000 + 100 x 10 = 1000 + 2000 + 1000 = 4000

2000 = 2 x 2000 + 4 x 4000 + 100 x 20 = 4000 + 16000 + 2000 = 22000

Real GDP = \sum(Base year (1999) price x Current year quantity)

1999 = 1 x 1000 + 1 x 2000 + 100 x 10 = 1000 + 2000 + 1000 = 4000

2000 = 1 x 2000 + 1 x 4000 + 100 x 20 = 2000 + 4000 + 2000 = 8000

GDP Deflator = (Nominal GDP / Real GDP) x 100

1999 = (4000/4000) x 100 = 100

2000 = (22000/8000) x 100 = 275

Inflation rate = % Change in GDP Deflator = (275/100) - 1 = 2.75 - = 1.75 = 175%

Therefore, CPI inflation > GDP Deflator inflation.